### v2

# Predicting the Performance of Graph Convolutional Networks with Spectral Properties of the Graph Laplacian

Abstract
A common observation in the Graph Convolutional Network (GCN) literature is that stacking GCN layers may or may not result in better performance on tasks like node classification and edge prediction. We have found empirically that a graph’s algebraic connectivity, which is known as the Fiedler value, is a good predictor of GCN performance. Intuitively, graphs with similar Fiedler values have analogous structural properties, suggesting that the same filters and hyperparameters may yield similar results when used with GCNs, and that transfer learning may be more effective between graphs with similar algebraic connectivity. We explore this theoretically and empirically with experiments on synthetic and real graph data, including the Cora, CiteSeer and Polblogs datasets. We explore multiple ways of aggregating the Fiedler value for connected components in the graphs to arrive at a value for the entire graph, and show that it can be used to predict GCN performance. We also present theoretical arguments as to why the Fiedler value is a good predictor.
Introduction
Many domains are naturally described by graphs, where nodes correspond to entities with properties and edges correspond to relationships among entities. Beyond the more obvious naturally occurring graphs, like social networks, citation networks, and molecular structures, one can represent text as an attributed graph (a dependency parse tree) and visual scenes as objects with properties that stand in spatial relationships to one another (Veličković 2023). Common inferential tasks performed on graphs include node prediction (e.g., there should be an object on which the cup is resting), link prediction (e.g., Alice probably knows Bob), and graph classification (e.g., that molecule is highly toxic).
Graph Convolutional Networks (GCNs) (Zhang et al. 2019) have emerged as a powerful tool for representation learning from graph data, with successful applications in physical systems modeling, molecular chemistry, biology, recommendation system design, traffic prediction, and other practical domains (Zhou et al. 2020). There are many forms of GCNs, but the basic idea is message passing. If each node initially contains a feature vector (though it is possible to have features on edges as well), those vectors are transmitted as messages to immediate neighbors and then all incoming messages are combined at a node to produce a new vector (Zhang et al. 2019). Adding layers to a GCN allows messages to spread farther in the graph. With $k$ layers, each node’s influence reaches all other nodes within $k$ hops. The resulting embeddings support node, edge, and graph classification tasks (Xu et al. 2023).
Despite their power and popularity, challenges remain with GCNs. They typically assume static graphs (Wu et al. 2022) and have difficulty scaling to large graphs (Mostafa and Nassar 2020). More relevant for this paper, unlike other forms of deep networks where increasing depth often leads to better representations, increasing depth with GCNs often has the opposite effect. For example, as the reach of nodes’ messages increases with GCN depth, the problem of over-smoothing arises, where the embeddings of all of the nodes become indistinguishable (Sun 2022). Similarly, over-squashing occurs as information from distant parts of the graph is squeezed (squashed) into a finite sized vector (Giraldo, Malliaros, and Bouwmans 2022). Adding layers also exacerbates problems with vanishing gradients (Li et al. 2019).
Collectively, this means that choosing the right number of layers for GCNs is a crucial design step, for which there is little guidance in the literature. In this paper we show that a graph’s algebraic connectivity (known as the Fiedler value or Fiedler eigenvalue), which is the first non-zero eigenvalue of the graph’s Laplacian matrix, can inform the choice of the number of GCN (Kipf and Welling 2016) layers.
Fisser et al.(Fesser and Weber 2025) introduce the Fiedler singular value of a gradient-derived graph to diagnose over-squashing by analyzing gradient flow bottlenecks in trained GNNs. In contrast, our work employs the Fiedler eigenvalue of the input graph Laplacian to derive a theoretical lower bound on feature variation resulting from Laplacian smoothing. While both approaches leverage spectral tools, our method is designed to provide insight into GCN behavior prior to training, enabling informed decisions about layer depth. This stands in contrast to Fisser’s post hoc diagnostic, which is applicable only after gradient-based learning has occurred.
Our contributions are as follows:
  • Introduction of algebraic connectivity of graphs as a way of characterizing how well GCNs will perform as a function of the number of convolutional layers.
  • Empirical demonstration of a “sweet spot” for the algebraic connectivity that indicates good GCN performance on node classification tasks on both synthetic and real-world data.
  • Theoretical arguments about why algebraic connectivity is a good predictor of GCN performance.
The remainder of this paper is organized as follows. The next section introduces concepts and notation relevant to graphs, algebraic connectivity, and GCNs. Following that is a theoretical exploration of algebraic connectivity, and then a series of experiments in which the algebraic connectivity of graphs is used to predict GCN performance.
Preliminaries
Let $G(V,E)$ be a graph $G$ where the set of vertices is $V$ and the set of edges is $E$. The adjacency matrix of $G$ is $A\in\{0,1\}^{n\times n}$. The degree matrix $D$ is defined as:
$D_{i,j}=\begin{cases}\text{degree}(v_{i}),&\text{if }i=j;\\ 0,&\text{otherwise.}\end{cases}$
The graph Laplacian matrix $L$ is defined as L=DAL=D-A (Grone, Merris, and Sunder 1990), which has an important role in defining convolution in GCNs (Daigavane, Ravindran, and Aggarwal 2021).
For any square matrix AA, it is the case that $Av=\lambda v$ where $v$ is an eigenvector of the matrix and $\lambda$ is the corresponding eigenvalue. Every square matrix has at least one eigenvalue (Mirchev 2016). In this work, we will be concerned with the eigenvalues of graph Laplacian matrices.
For any connected graph $G$ with $n$ vertices the Laplacian matrix of $G$ is symmetric and has real eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}....\leq\lambda_{n}$. The Fiedler value is the first nonzero eigenvalue of the Laplacian matrix $L$, i.e., $\lambda_{2}$ (Fiedler 1989). This value is associated with connectivity, among a few other graph-theoretic properties.
For a graph with $k$ disconnected components, the Laplacian matrix can be represented as a block diagonal matrix where each block corresponds to a single connected component of the graph. Let $L_{i}$ denote the graph Laplacian of the $i^{th}$ component. The spectrum of $L$ is defined by the union of the spectra of the $L_{i}$ (Von Luxburg 2007). Each $L_{i}$ contains eigenvalue 0 with multiplicity 1, with a corresponding eigenvector $\bf 1$}, so the number of 0 eigenvalues in $L$ equals the number of connected components in $G$. For any undirected graph, it is the case that algebraic connectivity is bounded from above by the vertex connectivity, or the number of nodes that must be removed to disconnect the graph (Barriere et al. 2013).
Convolutional neural networks (CNNs) applied to, for example, images use filters that compute weighted sums of values in a (typically) square region of adjacent pixels. GCNs use polynomial filters applied to the graph Laplacian matrix $L$. An example of a polynomial filter is shown below:
$S(L)=w_{0}I_{n}+w_{1}L+w_{2}L^{2}+...+w_{k}L^{K}=\sum\limits^{k}_{i=0}{w^{i}L^{i}}$
Here II is the identity matrix, the values of $w_{i}$ are the learned parameters of the polynomial filter, and $L^{i}$ denotes raising $L$ to the power $i$ or multiplying $i$ copies of $L$ together (Daigavane, Ravindran, and Aggarwal 2021). There are several types of polynomial filters, with the choice being domain or task specific (Gama et al. 2020).
One such filter is the Chebyshev polynomial, which follows De Moivre’s formula, meaning it can be generated recursively as follows: $P_{k}(x)=2xP_{k-1}(x)-P_{k-2}(x)$ where $P_{0}(x)=1$ and $P_{1}(x)=x$ (Defferrard, Bresson, and Vandergheynst 2016). They act as low pass filters in graph neural networks (He, Wei, and Wen 2022).
Among other Graph Neural Networks, GCNs use Chebyshev polynomials for convolution operations (Kipf and Welling 2016). GCN convolution layers aggregate information from the local neighborhood of nodes in a graph using a 1st-order approximation of the Chebyshev polynomial, aggregating information from 1-hop neighbors. For the layer-wise propagation rule, see Appendix A.2.
Although GCNs are an effective type of graph neural network for classification and node/link prediction, they suffer from over-smoothing (Elinas and Bonilla 2022). Adding multiple convolution layers causes node features to converge and the resulting model is unable to distinguish among them (Rusch, Bronstein, and Mishra 2023).
Algebraic Connectivity and GCNs
This section explores the connection between a graph’s Fiedler value (algebraic connectivity) and the operations used by GCNs, with the goal of understanding what the former may indicate about the latter.
Estimating the Topological Distance of a Graph
GCN models perform a type of Laplacian smoothing that combines local vertex data with graph topology in convolution layers. Vertices that are related topologically are likely to belong to the same class and smoothing these features increases their similarity, thereby aiding the classification process for GCNs (Li, Han, and Wu 2018). When nodes from different classes have small topological distances problems with over-smoothing are more likely to arise with deeper GCNs (Chen et al. 2020).
The mean distance is a measure of the topological distance of a graph. It is the average of all distances between distinct vertices of a graph G. Calculating the mean distance of a large and complex graph is computationally expensive, especially for real-world data. Approximating a bound on mean distance is often more practical. Topological distance plays a significant role in GCN performance, and the Fiedler value can be used to estimate the mean distance of a graph. It imposes both upper and lower bounds for a graph $G$ of order $n$ (Mohar 1991).
Let $\overline{m}$ denote the mean distance of a graph $G$ and let $\lambda_{2}$ denote the Fiedler value of the graph. Then a lower bound on the mean distance is:
$(n-1)\overline{m}(G)\geq\frac{2}{\lambda_{2}}+\frac{n-2}{2}.$
and an upper bound on the mean distance is:
$\overline{m}(G)\leq\frac{n}{n-1}\left(\left\lceil\frac{\Delta+\lambda_{2}}{4\lambda_{2}}\ln(n-1)\right\rceil+\frac{1}{2}\right).$
Here $\Delta$ is the maximum node degree of graph $G$.
From these bounds we can see that the Fiedler value and the mean distance of a graph have an inverse relationship. Higher mean distances mean, on average, that nodes are less closely connected and lower mean distances indicate closer connections. If the algebraic connectivity is high then the mean distance will be low and the graph will be more susceptible to over-smoothing as layers are added. High algebraic connectivity is also associated with expander-like graphs (see Appendix A.3).
A similar analysis (see the Appendix A.4) shows that the Fiedler value and the diameter of a graph are inversely related. If we apply too few GCN layers to a large-diameter graph, under‑reaching can occur. Conversely, large-diameter graphs are also more susceptible to over‑squashing, because exponentially many long‑range messages must be compressed into fixed‑size vectors as depth grows; small‑diameter graphs, in contrast, are more likely to encounter over‑smoothing. High range further exacerbates over‑squashing (Alon and Yahav 2020). If a graph’s diameter is large or small, as computed from the Fiedler value, we can adjust GCN depth accordingly.
Possible Optimum Feature Distance
It should be clear that the degree to which node features converge as a function of the number of rounds of graph convolution is central to the performance of GCNs. This is complicated by the fact that we’d like node features to converge differenly depending on the classes to which the nodes belong. What follows is a series of lemmas (with citations when they come from prior work) and a proposed theorem that will shed light on what the Fiedler value tells us about distances between feature vectors after Laplacian smoothing. we now establish a series of lemmas and conclude with a key theorem that connects the spectral properties of the graph, specifically the Fiedler value, to the distances between node features after Laplacian smoothing. The lemmas progressively build intuition, clarify important concepts like Dirichlet energy and feature distances, and lay the theoretical groundwork required for the main result. We begin with Lemma 1, a foundational result linking the Dirichlet energy to the total feature distance, providing a clear geometric interpretation of the Laplacian smoothing process.
Lemma 1.
Let $G=(V,E)$ be an undirected graph with degree matrix $D$, adjacency matrix AA, and Laplacian L=DAL=D-A. For each vertex $i$, let $v_{i}\in\mathbb{R}^{m}$ denote its feature vector, and let the feature matrix be $V=[v_{1}^{\top}\;\cdots\;v_{n}^{\top}]^{\top}\in\mathbb{R}^{n\times m}$. Then,
$\operatorname{Tr}(V^{\top}LV)=\sum_{\{i,j\}\in E}\|v_{i}-v_{j}\|_{2}^{2}.$
See Appendix A.5 for the proof.
Remark Lemma 1 can be viewed as a straightforward vector–valued generalisation of Marsden’s result (Marsden 2013, Lemma 3.6)
Interpretation.
The scalar $\operatorname{Tr}(V^{\top}LV)$ is often called the Dirichlet (or smoothness) energy of the graph signal VV(Giovanni et al. 2023). Lemma 1 shows that it is nothing more than the aggregate squared Euclidean distance between the features of adjacent vertices, thus providing a direct feature–distance interpretation. Note, however, that this quantity measures variation—it is not itself a graph-convolution operation; spectral filters and GCN layers are motivated by minimising this very score.
Lemma 2.
Let $G=(V,E)$ be an undirected graph with $|V|=n$, and let $L\in\mathbb{R}^{n\times n}$ denote its (combinatorial) Laplacian matrix. Then
$\lambda_{1}=0,\qquad u_{1}=\frac{1}{\sqrt{n}}\mathbf{1},$
where $\mathbf{1}=[1,\dots,1]^{\top}\in\mathbb{R}^{n}$, and $Lu_{1}=\lambda_{1}u_{1}$. This is a standard result in spectral graph theory (Chung 1997, Sec. 1.2).
See Appendix A.6 for the proof.
Lemma 3.
Let $G=(V,E)$ be a graph with $n$ vertices and Laplacian $L\in\mathbb{R}^{n\times n}$. Let $v_{i}\in\mathbb{R}^{m}$ be the feature vector of vertex $i$, and define $\bar{V}=[\,v_{1}^{\!\top}\;\dots\;v_{n}^{\!\top}]^{\top}\in\mathbb{R}^{n\times m}.$ Expand $\bar{V}$ in the orthonormal eigenbasis $\{u_{1},\dots,u_{n}\}\subset\mathbb{R}^{n}$ of $L$:
$\bar{V}\;=\;\sum_{i=1}^{n}u_{i}\,w_{i}^{\!\top},\qquad w_{i}\in\mathbb{R}^{m}.$
If the features are centered, i.e., $\sum_{i=1}^{n}v_{i}=0_{m}$, then $w_{1}=0_{m}$, and hence
$\bar{V}\;=\;\sum_{i=2}^{n}u_{i}\,w_{i}^{\!\top}.$
(See (Chung 1997, Chapter 1) for background on Laplacian eigenbases and orthogonality.)
Proof is provided in Appendix A.7.
Lemma 4.
Let $G$ be a graph with $n$ vertices and Laplacian $L\in\mathbb{R}^{n\times n}$. Stack the vertex features $v_{i}\in\mathbb{R}^{m}$ row-wise in the matrix $\bar{V}=[\,v_{1}^{\top}\dots v_{n}^{\top}]^{\top}\in\mathbb{R}^{n\times m}$. Suppose that
$\bar{V}\;=\;\sum_{i=1}^{n}\bar{u}_{i}\,w_{i}^{\!\top},\qquad w_{i}\in\mathbb{R}^{m},$
where $\{\bar{u}_{i}\}_{i=1}^{n}\subset\mathbb{R}^{n}$ is an orthonormal eigenbasis of $L$. If the features are centered, $\sum_{i=1}^{n}v_{i}=0_{m}$, then
$\sum_{i=1}^{n}\|v_{i}\|_{2}^{2}\;=\;\sum_{i=2}^{n}\|w_{i}\|_{2}^{2}.$
See Appendix A.8 for the proof.
Remark. Lemma 4 can be viewed as a matrix-valued extension of Parseval’s identity in the Laplacian spectral domain. It relies on the orthonormality of the Laplacian eigenbasis and the fact that centering removes the component along the constant eigenvector. Analogous scalar results are well established in spectral graph theory and graph signal processing (Chung 1997; Shuman et al. 2013).
Lemma 5.
Let $G$ be a graph with $n$ vertices, Laplacian $L\in\mathbb{R}^{n\times n}$ and eigen-pairs $\{(\lambda_{i},u_{i})\}_{i=1}^{n}$ ordered $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$. Stack the vertex features $v_{i}\in\mathbb{R}^{m}$ row-wise in the matrix $\bar{V}=[\,v_{1}^{\top}\dots v_{n}^{\top}]^{\top}\in\mathbb{R}^{n\times m}$ and expand it in that eigenbasis
$\bar{V}=\sum_{i=1}^{n}u_{i}\,w_{i}^{\!\top},\qquad w_{i}\in\mathbb{R}^{m}.$
If the features are centered $\bigl{(}\sum_{i=1}^{n}v_{i}=0_{m}\bigr{)}$, then $w_{1}=0_{m}$ and
$\sum_{(i,j)\in E}\!\|v_{i}-v_{j}\|_{2}^{2}=\operatorname{Tr}(\bar{V}^{\top}L\bar{V})=\sum_{i=2}^{n}\lambda_{i}\,\|w_{i}\|_{2}^{2}.$
The proof is provided in Appendix A.9.
Remark. Lemma 5 extends the classical spectral decomposition of the Laplacian quadratic form to matrix-valued signals. In the scalar case, this identity expresses the Dirichlet energy as a weighted sum of squared spectral coefficients, with weights given by the Laplacian eigenvalues. See (Chung 1997, Chapter 1) and (Shuman et al. 2013) for foundational treatments in spectral graph theory and graph signal processing.
Theorem 1 (Fiedler Bound for Minimal Dirichlet Energy).
Let $G=(V,E)$ have Laplacian $L\in\mathbb{R}^{n\times n}$ with eigenvalues $0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}$ (the eigenvalue $\lambda_{2}$ is the Fiedler value). For vertex features $v_{i}\in\mathbb{R}^{m}$ satisfying
$\sum_{i=1}^{n}\|v_{i}\|_{2}^{2}=n,\qquad\sum_{i=1}^{n}v_{i}=0_{m},$
we always have
$\sum_{(i,j)\in E}\|v_{i}-v_{j}\|_{2}^{2}\;\;\geq\;n\,\lambda_{2},$
and the lower bound $n\lambda_{2}$ (the Fiedler bound) is attained iff all energy lies in the Fiedler mode, i.e.
$w_{2}=\sqrt{n}\,q,\ q\in\mathbb{R}^{m},\qquad w_{3}=\dots=w_{n}=0_{m},$
where the $w_{i}$ are the coefficients in the expansion $\bar{V}=\sum_{i=2}^{n}u_{i}w_{i}^{\!\top}$.
See Appendix A.10 for the proof of Theorem 1. It is important to note that the proposed theorem establishes a connection between the minimum feature distance and the Fiedler value, and comparing the actual feature distances with this minimum can give insight into the effects of adding GCN layers. If the empirical feature distribution is close to the minimum distance for a particular graph, then the graph features are smooth and GCN performance classification tasks should be good. On the other hand, adding layers to the GCN quickly lead to over-smoothing.
While the variational characterization of the Fiedler value via the Rayleigh quotient is well known in spectral graph theory, our contribution lies in its reinterpretation as a lower bound on the feature variation achievable through Laplacian smoothing. We explicitly frame this as a constraint for GCN feature evolution and propose a normalized diagnostic score $(\rho_{k})$, which allows one to monitor how close a representation is to this theoretical limit. To our knowledge, this is the first application of the Fiedler bound as a dynamic smoothing indicator for GCN layer selection.
Feature–distance score.
Theorem 1 shows that $n\lambda_{2}$—the Fiedler bound—is the smallest total feature distance any centered, unit-energy signal can achieve on the graph, so we use $n\lambda_{2}$ as a gold standard for smoothness. For the hidden representation $V^{(k)}$ produced after the $k$-th GCN layer, define
$\rho_{k}\;=\;\frac{\operatorname{Tr}\!\bigl{(}(V^{(k)})^{\top}LV^{(k)}\bigr{)}}{n\lambda_{2}}.$ (1)
Before training ($k=0$). With $V^{(0)}\!=V_{\text{in}}$ we have
$\rho_{0}\;=\;\frac{\operatorname{Tr}\!\bigl{(}V_{\text{in}}^{\top}LV_{\text{in}}\bigr{)}}{n\lambda_{2}}.$ (2)
A large $\rho_{0}$ (say $\rho_{0}\!\gg\!1$) indicates that the raw features vary sharply across edges, leaving the network ample room to smooth them.
If $\rho_{k}$ stays $\gg 1$ after a few layers, additional depth can legitimately reduce feature variation; when $\rho_{k}\!\approx\!1$, further layers are likely to push the model into the over-smoothing regime.
Experiments and Results
This section describes experiments that seek to establish empirical evidence for the theoretical connection between the Fieldler value and GCN performance described in the previous section. In our experiments, we implemented the Graph Convolutional Network (GCN) architecture originally proposed by Kipf et al. (Kipf and Welling 2016), closely following their formulation and training methodology. Our current formulation assumes that label similarity is correlated with topological closeness (homophily), which underlies the use of Laplacian smoothing.
We evaluate our method on two real-world citation network datasets Cora and CiteSeer (Yang, Cohen, and Salakhudinov 2016)—as well as the PolBlogs dataset, originally introduced by Adamic and Glance (Adamic and Glance 2005) and accessed via the Network Repository (Rossi and Ahmed 2015). We focus on the citation networks Cora and Citeseer, the same benchmarks used in several recent over‑smoothing studies e.g. the theoretical analysis of mean‑aggregation smoothing by Keriven at el (Keriven 2022).
Additionally, we also used synthetic datasets. Nodes in the synthetic datasets belong to one of three classes, with the node features being drawn from one of three normal distributions, each paired with a class. The means and variances of those normals are as follows: (250, 50), (100, 90), and (400, 200). Synthetic graphs enable controlled variation of structural parameters such as edge density and connectedness which in turn affect the spectral properties of the Laplacian. Unlike real-world graphs, where many factors (e.g., label distribution, feature heterogeneity) may co-vary, these graphs allow us to isolate the specific influence of algebraic connectivity on GCN performance. Real-world graphs can have multiple connected components. The Fielder value, $\lambda_{2}$, in such cases is 0. To addrress this we compute $\lambda_{2}$ for each component and aggregate them using a weighted average, where the weights are the ratios of the number of nodes in each component to the total number of nodes in the graph. In this way, larger connected components have greater influence on the aggregate Fiedler value that will be used to make predictions about GCN performance. The number of nodes, edges, and features in the synthetic datasets were varied, and the values of $\lambda_{2}$ were calculated. Then GCNs were trained with up to five layers for node classification.
Figure 1: Accuracy of GCN layers for four graph datasets in various conditions.
Figure 1(a) demonstrates the node classification accuracy of Graph Convolutional Networks (GCNs) across graphs with varying edge densities. Specifically, four graphs, each consisting of 500 nodes and 100-dimensional node features but differing in the number of edges, are analyzed. The horizontal axis represents the number of GCN layers, and the vertical axis indicates the classification accuracy. In particular, regardless of edge count, three-layer GCNs consistently achieve the highest accuracy. In Figure 1(a) the graphs are arranged along the horizontal axis in descending order according to their edge counts illustrating how increased connectivity impacts these metrics. Fiedler values, again ordered by edge count.
Figure 2: Structural metrics derived from the four 500‑node graphs used in Fig. 1(a). (a) Average of Fiedler values (b) Weighted average of Fiedler values.
In Figure 2(a), a comparison is depicted among the graphs of Figure 1(a) based on the average of the Fiedler values, total components of the graph, and total number of connected subgraphs. Similarly, Figure 2(b) further refines this analysis by considering the weighted average of Fiedler values, from Figure 1(a) in descending order based on the number of edges in the graph. The weighted average calculation involves multiplying each component’s Fiedler value by its corresponding number of nodes, summing these products, and then dividing by the total number of nodes. This metric captures a nuanced perspective of the graph’s spectral connectivity relative to its structural complexity.
Figure 3: Structural metrics derived from the four 1000‑node graphs used in Fig. 1(b). (a) Average of Fiedler values (b) Weighted average of Fiedler values.
Figure 3(a) depicts the total number of components, total number of connected subgraphs, and an average of Fiedler values of Figure 1(b). The data of the graphs are also presented in descending order based on the number of edges, highlighting how these structural attributes vary systematically with graph connectivity. In the same way, Figure 3(b) further elaborates by displaying the weighted average Fiedler values of Figure 1(b). This metric emphasizes how spectral connectivity scales with the structural composition of the graph.
Figure 4: Structural metrics derived from the four 1500‑node graphs used in Figure 1(c). (a) Average of Fiedler values (b) Weighted average of Fiedler values.
Figure 1(c) illustrates the accuracy of GCN for four graph datasets of 1500 nodes and 100 node features. Classification accuracy decreases at the extremes of edge connectivity, being lower for both densely and sparsely connected graphs.
Figure 4(a) illustrates the average of Fiedler values, number of total components, and connected subgraphs of Figure 1(c). A comparison is made among the four graphs in descending order based on the number of edges. Similarly, Figure 4(b) represents the weighted average of the Fiedler values of Figure 1(c).
Figure 5: Structural metrics derived from the four 1000‑node graphs used in Fig. 1(d). (a) Average of Fiedler values (b) Weighted average of Fiedler values.
Figure 1(d) presents node classification accuracy across multiple GCN layers on a single initial graph comprising 1000 nodes, 100-dimensional node features, and initially 950 edges. To investigate how connectivity influences accuracy and spectral properties, edges were gradually and randomly removed (from 950 down to 500, 250, and 100). The horizontal axis denotes the number of GCN layers, while the vertical axis indicates accuracy. Consistent with previous observations, accuracy peaks at three layers regardless of the graph’s connectivity, highlighting an optimal GCN depth under these conditions. The data are organized in descending order by edge count, clearly showing how connectivity reduction affects these structural measures.
Figure 5(a) and Figure 5(b) depict the average of Fiedler values and also the weighted average of the Fiedler values of Figure 1(d). Comparison is made in descending order based on the number of edges. This weighted metric further clarifies how decreasing graph connectivity directly affects spectral characteristics critical to GCN performance. Overall, these figures collectively emphasize how graph structural properties, particularly as captured by Fiedler values and connectivity metrics, correlate with GCN performance and optimal model depth.
Table 1 depicts the performance of GCN in Cora CiteSeer and Polblogs datasets. The accuracy of applying GCN is dropped immediately after applying 3 convolution layers for Cora and CiteSeer dataset and for Polblogs dataset it takes 5 layers. We have preprocessed the PolBlogs dataset by removing communities that have only one member to successfully perform the node classification task. After preprocessing, the number of nodes is 1224 and the classes are 11. Since the dataset does not include node features, we assign each node a unique one-hot encoded vector as its feature representation, resulting in an identity matrix of size $N\times N$, where NN is the number of nodes in the preprocessed graph.
Table 2 illustrates the average Fiedler values and weighted average Fiedler values of Cora, CiteSeer and Polblogs datasets.
Dataset Nodes Feat. Edges Acc@2 Acc@3 Acc@4 Acc@5
Cora 2,708 1,433 10,556 0.8150 0.7910 0.7410 0.6730
CiteSeer 3,327 3,703 9,104 0.7120 0.6640 0.6030
Polblogs 1,490 1,224 16,718 0.9565 0.9592 0.9620 0.9484
Table 1: Accuracy of GCN with varying convolution layers
Dataset Avg. of FV Weighted avg. of FV
Cora 1.62833 0.1024
CiteSeer 1.54106 0.4205
Polblogs 1.50001 0.04615
Table 2: Average of Fiedler values and weighted average of Fiedler values of real-world datasets
Our experiments reveal an optimal range of Fiedler values associated with strong GCN performance. Specifically, we observe diminished classification performance when algebraic connectivity is either too low or too high, even with identical numbers of nodes and features. Empirically, average Fiedler values exceeding approximately 1.4 consistently correlate with reduced performance. In contrast, we find the strongest performance when the weighted average of Fiedler values falls between approximately 0.5 and 1.2. This trend holds consistently across multiple experiments, indicating suboptimal performance for both excessively dense (high edge count) and excessively sparse (low edge count) graph structures. This enables the use of $\lambda 2$ as a pre-training diagnostic to inform architecture choices and potentially reduce expensive hyperparameter search
Conclusion
Graph Convolutional Networks (GCNs) face well-documented performance limitations, particularly in relation to the structural properties of their input graphs. This work highlights the Fiedler value as a theoretically grounded and empirically informative metric for assessing and predicting GCN performance. By quantifying algebraic connectivity, the Fiedler value offers practical guidance on whether a given graph falls within a range where shallow or deep GCNs—or GCNs at all—are likely to be effective. Our findings suggest that incorporating spectral characteristics, especially the Fiedler value, into model selection and experimental design can improve performance across diverse datasets.
The broader impacts of this work include reducing the time to create effective graph algorithms by choosing an appropriate number of convolution layers quickly, resulting in fewer training runs and a smaller carbon footprint. This study is limited to node classification tasks on graphs with multiple components. Future work will explore broader empirical validation, including alternative tasks and metrics such as feature distance scores, to further elucidate the role of spectral properties in GCN performance.
A. Appendix
A.1 Notation
Notations are depicted in Table 3
Symbol Meaning Shape
$\mathcal{V}$ vertex set
$n:=|\mathcal{V}|$ number of vertices
$\mathbf{A}$ adjacency matrix $n\times n$
$\mathbf{D}$ degree matrix $n\times n$
$\mathbf{L}=\mathbf{D}-\mathbf{A}$ combinatorial Laplacian $n\times n$
$\mathbf{X}=[x_{1}^{\top}\,\dots\,x_{n}^{\top}]^{\top}$ input feature matrix $n\times m$
$V^{(k)}$ hidden feature matrix after $k$-th GCN layer $n\times m_{k}$
$u_{i},\lambda_{i}$ $i$-th eigenpair of $\mathbf{L}$ $n$, scalar
$w_{i}$ spectral coefficient vectors $m$
$\delta_{ij}$ Kronecker delta (=1=1 if $i=j$)
$\rho_{k}$ normalised feature–distance score (Eq. 1) scalar
Table 3: Glossary of symbols used throughout the paper.
A.2 Layer-wise Propagation rule of GCN
$H^{(l+1)}=\sigma\left(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right)$ (3)
where:
  • $H^{(l)}$ is the feature matrix of layer $l$
  • $W^{(l)}$ is the learnable weight matrix for layer $l$
  • $\tilde{A}=A+I_{N}$ is the adjacency matrix with added self-loop
  • $\tilde{D}$ is the degree matrix corresponding to $\tilde{A}$
  • $\sigma$ is a non-linear activation function
$H^{(0)}$ is the input feature matrix, and the output of the final layer $H^{(L)}$ (after $L$ layers) is used for tasks like classification. In the above formula, $\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}$ is derived from the normalized graph Laplacian (Kipf and Welling 2016).
A.3 Predicting Expander-like Graph
A larger Fiedler value indicates an expander-like graph (Kohayakawa, Rödl, and Schacht 2003), and GCNs are known to suffer from information loss in such graphs (Oono and Suzuki 2019). For an arbitrary vertex $\omega\in V(G)$ and a non-negative whole number $k$, let $R_{k}=R_{k}(\omega)$ be the set of vertices of $G$ which are at distance at most $l$ from $\omega$ and $r_{k}=|R_{k}|$. $r_{k}$ has exponential growth, i.e. $r_{k}\geq\alpha^{k}$ where $\alpha$ is a constant bounded below as a function of the Fiedler value ($\lambda_{2}$) (Mohar 1991). Larger Fiedler values indicate that exponentially more information reaches nodes with additional rounds of convolution, increasing the chances of over-squashing.
A.4 Under-Reaching and Over-Squashing
The diameter of a graph is the longest shortest path between any two nodes in the graph, and it has a profound effect on GCN performance. Computing the exact diameter of a large graph is computationally intensive and often impractical, making diameter approximation a more feasible approach. The Fiedler value can be used to estimate the diameter of a real‑world graph. This assessment may assist in mitigating issues related to under‑reaching or over‑squashing. When a node does not receive needed information about other nodes that are far away, this is called under‑reaching (Alon and Yahav 2020). Recall that if we apply $l$ convolution layers in a GCN, nodes that are separated by a shortest path of more than $l$ hops will not be “aware” of each other. On the other hand, GCNs suffer from over‑squashing when exponentially growing information about increasingly distal parts of the graph is forced into a fixed‑size vector. GCN models applied to graphs with large diameters often face over‑squashing (Alon and Yahav 2020). In both cases, knowing the diameter of a graph is useful to determine whether under‑reaching or over‑squashing might be a problem. Again, the Fiedler value can be used to bound the diameter of a graph $G$ of order nn(Mohar 1991).
The lower bound on the diameter is:
$\operatorname{diam}(G)\;\geq\;\frac{4}{n\lambda_{2}}.$
and the upper bound is:
$\operatorname{diam}(G)\;\leq\;2\Bigl{\lceil}\sqrt{\frac{2\Delta}{\lambda_{2}}}\,\log_{2}n\Bigr{\rceil}.$
Here $\Delta$ is the maximum node degree of graph $G$.
A.5 Proof of Lemma 1
Proof.
Using the identity L=DAL=D-A and the cyclic property of the trace
$\operatorname{Tr}(V^{\top}LV)=\operatorname{Tr}(V^{\top}DV)-\operatorname{Tr}(V^{\top}AV).$
Since the graph is undirected, each unordered edge $\{i,j\}\in E$ appears exactly once, and $D_{ii}=d_{i}=\sum_{j}A_{ij}$. Therefore,
$\operatorname{Tr}(V^{\top}DV)=\sum_{i=1}^{n}d_{i}\|v_{i}\|_{2}^{2}=\sum_{\{i,j\}\in E}\left(\|v_{i}\|_{2}^{2}+\|v_{j}\|_{2}^{2}\right),$
and
$\operatorname{Tr}(V^{\top}AV)=\sum_{i,j}A_{ij}v_{i}^{\top}v_{j}=2\sum_{\{i,j\}\in E}v_{i}^{\top}v_{j}.$
Subtracting the two expressions gives:
$\operatorname{Tr}(V^{\top}LV)=\sum_{\{i,j\}\in E}\left(\|v_{i}\|_{2}^{2}+\|v_{j}\|_{2}^{2}-2v_{i}^{\top}v_{j}\right)\\ =\sum_{\{i,j\}\in E}\|v_{i}-v_{j}\|_{2}^{2},$
which completes the proof. ∎
A.6 Proof of Lemma 2
Proof.
By definition, L=DAL=D-A, where $D$ is the degree matrix and AA is the adjacency matrix. Since $G$ is undirected, each row of $L$ sums to zero. That is, for all $i\in\{1,\dots,n\}$, we have $\sum_{j=1}^{n}L_{ij}=0$. It follows that $L\mathbf{1}=0$, so $\lambda_{1}=0$ is an eigenvalue of $L$, with eigenvector $\mathbf{1}$. Normalizing $\mathbf{1}$ yields the unit eigenvector $u_{1}=\tfrac{1}{\sqrt{n}}\mathbf{1}$. ∎
A.7 Proof of Lemma 3
Proof.
The first Laplacian eigenvector is the normalized all-ones vector $u_{1}=\tfrac{1}{\sqrt{n}}\mathbf{1}$ (see Lemma 2). Left-multiplying $(\ast)$ by $\mathbf{1}^{\top}$ gives
$\mathbf{1}^{\top}\bar{V}\;=\;\sum_{i=1}^{n}\mathbf{1}^{\top}u_{i}\,w_{i}^{\!\top}\;=\;\sqrt{n}\,w_{1}^{\!\top},$
since $\mathbf{1}^{\top}u_{i}=0$ for $i\geq 2$ by orthogonality. The centering condition implies $\mathbf{1}^{\top}\bar{V}=0$, so $w_{1}^{\!\top}=0$ and thus $w_{1}=0_{m}$. ∎
A.8 Proof of Lemma 4
Proof.
Centering implies $\mathbf{1}^{\top}\bar{V}=0_{1\times m}$; since $\mathbf{1}^{\top}\bar{u}_{i}=0$ for $i\geq 2$ and $\mathbf{1}^{\top}\bar{u}_{1}=\sqrt{n}$, we must have $w_{1}=0_{m}$. Using the orthogonality $\bar{u}_{i}^{\top}\bar{u}_{j}=\delta_{ij}$,
$\operatorname{Tr}(\bar{V}^{\top}\bar{V})$ $=$ $\operatorname{Tr}\!\Bigl{(}\bigl{[}\sum_{i=2}^{n}w_{i}\,\bar{u}_{i}^{\top}\bigr{]}\bigl{[}\sum_{j=2}^{n}\bar{u}_{j}\,w_{j}^{\top}\bigr{]}\Bigr{)}$
$=$ $\sum_{i=2}^{n}\operatorname{Tr}(w_{i}w_{i}^{\top})$
$=$ $\sum_{i=2}^{n}\|w_{i}\|_{2}^{2}.$
But $\operatorname{Tr}(\bar{V}^{\top}\bar{V})=\sum_{i=1}^{n}\|v_{i}\|_{2}^{2}$, completing the proof. ∎
A.9 Proof of Lemma 5
Proof.
Centering forces $w_{1}=0_{m}$ exactly as in Lemma 3; thus $\bar{V}=\sum_{i=2}^{n}u_{i}w_{i}^{\!\top}$. Using cyclicity of the trace,
$\operatorname{Tr}(\bar{V}^{\top}L\bar{V})$ $=$ $\operatorname{Tr}\!\Bigl{(}\Bigl{[}\sum_{i=2}^{n}w_{i}\,u_{i}^{\top}\Bigr{]}L\Bigl{[}\sum_{j=2}^{n}u_{j}\,w_{j}^{\!\top}\Bigr{]}\Bigr{)}$
$=$ $\sum_{i=2}^{n}\sum_{j=2}^{n}w_{i}^{\top}\underbrace{u_{i}^{\top}Lu_{j}}_{=\;\lambda_{j}\delta_{ij}}w_{j}$
$=$ $\sum_{i=2}^{n}\lambda_{i}\,w_{i}^{\top}w_{i}$
$=$ $\sum_{i=2}^{n}\lambda_{i}\,\|w_{i}\|_{2}^{2}.$
The first equality uses the identity $\sum_{(i,j)\in E}\|v_{i}-v_{j}\|_{2}^{2}=\operatorname{Tr}(\bar{V}^{\top}L\bar{V})$ from Lemma 1. ∎
A.10 Proof of Theorem 1
Proof of Theorem 1.
Step 1 : Rayleigh–Ritz expression for the objective. By Lemma 5 the total feature distance can be written as the quadratic form
$\mathcal{E}(\bar{V})\;=\;\sum_{(i,j)\in E}\!\|v_{i}-v_{j}\|_{2}^{2}\;=\;\sum_{i=2}^{n}\lambda_{i}\,\|w_{i}\|_{2}^{2},$
where the coefficient vectors $w_{i}\in\mathbb{R}^{m}$ come from the outer‑product expansion $\bar{V}=\sum_{i\geq 2}u_{i}w_{i}^{\!\top}$.
Step2: Constraint in the coefficient space. Lemma 4 (with $w_{1}=0_{m}$, enforced by the centering assumption) converts the Frobenius‑norm constraint into
$\sum_{i=2}^{n}\|w_{i}\|_{2}^{2}\;=\;n.$
Step3: Optimisation. Because the eigenvalues are ordered $0=\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\dots\leq\lambda_{n}$ and the terms $\|w_{i}\|_{2}^{2}$ are non‑negative,the Rayleigh–Ritz energy $\mathcal{E}(\bar{V})=\sum_{i=2}^{n}\lambda_{i}\|w_{i}\|_{2}^{2}$ is minimised, subject to the norm constraint $\sum_{i=2}^{n}\|w_{i}\|_{2}^{2}=n$, by allocating the entire norm budget $n$ to the smallest eigenvalue
$\lambda_{2}$:
$\|w_{2}\|_{2}^{2}=n,\qquad\|w_{i}\|_{2}^{2}=0\ \ (i\geq 3).$
Setting $w_{2}=\sqrt{n}\,q$ with any unit vector $q\in\mathbb{R}^{m}$ realises that allocation and yields
$\mathcal{E}(\bar{V}_{\min})=\lambda_{2}\,\|w_{2}\|_{2}^{2}=n\,\lambda_{2}.$
Step4: Uniqueness of the minimiser. If a feasible set $\{w_{i}\}_{i=2}^{n}$ achieves the same value $n\lambda_{2}$, then necessarily $\sum_{i=2}^{n}(\lambda_{i}-\lambda_{2})\|w_{i}\|_{2}^{2}=0$; because each summand is non‑negative and $\lambda_{i}>\lambda_{2}$ for $i\geq 3$, this forces $\|w_{i}\|_{2}^{2}=0$ for all $i\geq 3$. Hence every minimiser must indeed place all energy in the Fiedler mode. ∎